OpenStudy (anonymous):

convert 320 degrees to radians

OpenStudy (amistre64):

degrees need pi; so multiply it by pi and divide it by 180

OpenStudy (lalaly):

320 *pi /180

OpenStudy (anonymous):

\[320\times \pi \div180\]=\[2\pi\]